HIGH RESOLUTION INFRARED SPECTROSCOPY IN P ARA-H_{2} CRYSTALS

\underline{TAKESHI OKA}, Department of Chemistry, The University of Chicago, Chicago, IL 60637.

Because of the weak intermolecular interactions, molecular rotations and vibrations are nearly free in para-H_{2} crystals, and their quantum states have long relaxation times^{a}. Vibration-rotation transitions appear extremely sharp with line widths that are smaller than those of gaseous lines by one to two orders of magnitude (Δν/ν ∼ 10^{−7} − 10^{−8}). We can thus conduct high resolution spectroscopy and study solid states with unprecedented accuracy and clarity. I will discuss pure rotational (e.g. J = 6 ← 0), pure vibrational (v = 1 ← 0, J = 0 ← 0), and vibration-rotation pair transitions with emphasis on qualitative and quantitative difference between solid and gaseous spectroscopies.

aT. Oka, Ann. Rev. Phys. Chem. \underline{44}, 299 (1993).